OpenStudy (anonymous):

a line has a slope equal to 1, amd it passes through the point (-2,-2). What is its equation?

OpenStudy (texaschic101):

y - y1 = m(x - x1) y - (-2) = 1(x - (-2) y + 2 = x + 2 y = x + 2 - 2 y = x + 0

OpenStudy (anonymous):

THANK YOUUU

OpenStudy (texaschic101):

no problem :)